### v1

# Evaluating the Elementary Multilingual Capabilities of Large Language Models with MultiQ

Abstract
Large language models (LLMs) need to serve everyone, including a global majority of non-English speakers. However, most LLMs today, and open LLMs in particular, are often intended for use in just English (e.g. Llama2, Mistral) or a small handful of high-resource languages (e.g. Mixtral, Qwen). Recent research shows that, despite limits in their intended use, people prompt LLMs in many different languages. Therefore, in this paper, we investigate the basic multilingual capabilities of state-of-the-art open LLMs beyond their intended use. For this purpose, we introduce MultiQ, a new silver standard benchmark for basic open-ended question answering with 27.4k test questions across a typologically diverse set of 137 languages. With MultiQ, we evaluate language fidelity, i.e. whether models respond in the prompted language, and question answering accuracy. All LLMs we test respond faithfully and/or accurately for at least some languages beyond their intended use. Most models are more accurate when they respond faithfully. However, differences across models are large, and there is a long tail of languages where models are neither accurate nor faithful. We explore differences in tokenization as a potential explanation for our findings, identifying possible correlations that warrant further investigation.
Evaluating the Elementary Multilingual Capabilities
of Large Language Models with
MultiQ
Carolin Holtermann1*, Paul Röttger2*, Timm Dill1, Anne Lauscher1 1Data Science Group, University of Hamburg, Germany 2Bocconi University, Italy carolin.holtermann@uni-hamburg.de
**footnotetext: Equal Contribution
1 Introduction
Languages other than English remain underrepresented and underserved by state-of-the-art language technologies, posing a barrier to equal and inclusive AI (Bender, 2011; Joshi et al., 2020). While proprietary large language models (LLMs) like GPT-4 (OpenAI, 2023) may answer questions and follow instructions in many different languages, even the best and most popular open LLMs are much more restricted in their language coverage: Llama2-chat (Touvron et al., 2023b), for example, is “intended for commercial and research use in English”. [^1] https://huggingface.co/meta-llama/Llama-2-7b-hf Yi is “bilingual” in English and Chinese [^2] https://huggingface.co/01-ai/Yi-34B-Chat, and Mistral-7b-instruct (Jiang et al., 2023) “only works in English”. [^3] https://mistral.ai/news/la-plateforme/
Figure 1: The 137 languages covered in our MultiQ question dataset. We show their geographic location according to the WALS database and indicate their corresponding language family through colors.
Even though most open LLMs are restricted in their intended use to one or a handful of languages, datasets of real-world LLM usage show that people prompt LLMs in many different languages, often beyond their intended use (Zheng et al., 2023; Ouyang et al., 2023; Zhao et al., 2024). This has motivated initial research into the multilingual capabilities of monolingual models (Armengol-Estapé et al., 2022; Lai et al., 2023a). However, this research has mostly focused on older proprietary LLMs and on a relatively small number of languages and/or specific tasks.
In this paper, we investigate the basic multilingual capabilities of a variety of state-of-the-art chat-optimized open LLMs across a typologically diverse set of 137 languages. Specifically, we ask two main research questions that correspond to two dimensions of capability: 1) What is the multilingual language fidelity of current chat-optimized open LLMs, and 2) What is the multilingual question answering (QA) accuracy of current chat-optimised open LLMs? Language fidelity describes the ability to respond to prompts in the prompted language. QA accuracy describes the ability to give correct answers to open-ended questions, in the prompted language or any other. An ideal multilingual model would give answers that are both faithful and correct.
To answer our two research questions, we introduce MultiQ, a new silver standard benchmark for basic open-ended question answering comprising 27,400 test prompts across 137 typologically diverse languages. We create MultiQ by compiling 200 English questions that are simple yet realistic and diverse, and translating them automatically to 136 other languages. We evaluate QA accuracy on MultiQ using a GPT-4 classifier and language fidelity using GlotLID (Kargaran et al., 2023). MultiQ is a silver standard because automated translation and evaluation introduce some noise into the results. However, we validate through expert annotation that this noise is likely small, thus demonstrating that MultiQ can provide valuable evidence on basic multilingual capabilities. Concretely, we use MultiQ to make four main findings:
  1. 1.
    Language Fidelity: While some open models (e.g. Llama2) mostly respond in English regardless of the input language, other models (e.g. Mistral) respond faithfully despite their intended use being monolingual.
  2. 2.
    QA Accuracy: On MultiQ, all models tend to perform best in English, with some performing similarly well in up to 20 other languages (e.g. Mixtral). Across models, there is a long tail of languages with very poor accuracy.
  3. 3.
    Interaction: Increased language fidelity may positively impact answer accuracy, given that model answers in the same language as prompted tend to be more accurate.
  4. 4.
    Explanation: With regard to the tokenization strategy of models for different languages, models tend to achieve higher accuracy on languages they can tokenize into subwords instead of characters or ASCII tokens.
2 The MultiQ Dataset
MultiQ is a collection of 27,400 simple open-ended questions across 137 typologically diverse languages. The questions cover different topics ranging from algebra to geography to astronomy. Questions in each language are parallel to each other. The open-ended question format is consistent with real-world LLM usage. Additionally, the open-endedness minimizes the likelihood of correct answers given by chance.
2.1 Dataset Creation
We created MultiQ in two steps. First, we compiled an initial set of English questions. Second, we automatically translated these prompts into 136 typologically diverse languages.
For the initial English questions, we used two different sources to increase question diversity. a) We collected 100 questions from the LMSYS-Chat-1M dataset (Zheng et al., 2023), which catalogs real-world user interactions with LLMs. Specifically, we sampled all single-sentence English-language and sorted them by frequency. Then, we manually selected from the top until we reached 100 questions. This portion of our data directly reflects real-world LLM usage. b) We manually created another set of 100 questions evenly spread across 10 different subjects at elementary to middle school level (e.g.   mathematics and geography). To maximize the diversity of the questions, we prompted GPT-4 to provide us with a set of simple and clear questions with simple and clear answers for each of the subjects. We then iterated and manually selected questions until we reached 10 questions per subject. This portion of our data expands MultiQ’s topical coverage.
In both the LMSYS and the GPT-4 portions of our data, we manually selected only questions that are simple, factual, and target common knowledge. This is because with MultiQ we want to test basic multilingual capabilities, not complex reasoning. Questions must also have unambiguous answers that are culturally and temporally invariant. This is to minimize discrepancies introduced by translation as well as temporal degradation of our dataset. Table 1 shows English example prompts.
Source Example Prompt
GPT-4 What is the chemical formula for water?
Who was the first Emperor of China?
What is a galaxy?
LMSYS Was the year 2000 a leap year?
What is 2 + 2 * 3?
How many feet does a chicken have?
Table 1: Examples of English questions covered in MultiQ. We present three prompts from each source (GPT-4 and LMSYS) covering different domains.
To translate our English questions into other languages, we used the v3 Google Translate API. [^4] https://translation.googleapis.com/v3/ Specifically, we translated the 200 initial English questions into all 136 other languages covered by the API as of February 2024, resulting in a total of 27,400 parallel questions in MultiQ. The decision to use automated translation was driven by the constraints of our research budget, which made manual translation infeasible. The benefit of automated translation is that we can cover many more languages. Next, we discuss how we validated the quality of the translations, and demonstrate the typological diversity of the 137 languages we cover.
2.2 Validation of Translations
We asked native speakers to annotate the correctness of the 200 translated MultiQ questions for 10 languages: Chinese, French, Farsi, German, Indonesian, Italian, Japanese, Russian, Spanish, and Urdu. Across these languages, annotators marked an average of 94.4% of translations as correct. Translations were least accurate for Indonesian and Farsi, at 89.0%, while the translations for Italian and German were the most accurate, at 99.0%. We present the full results in the Appendix A. Qualitatively, several annotators stated that some translations, while accurate in content, tended to be literal, word-for-word translations rather than natural expressions. Overall, the automated translation introduces some noise into MultiQ, but our validation results suggest that the amount of noise is limited. This is why we frame MultiQ as a silver standard benchmark that can provide meaningful insights into basic multilingual capabilities, even if exact results on individual test cases may not be perfectly reliable.
2.3 Typological Diversity
MultiQ covers a total of 137 languages. To demonstrate their typological diversity, we follow best practices suggested by Ploeger et al. (2024), analyzing both between-language distances as well as overall typological feature coverage.
To estimate language distances, we use the lang2vec toolkit (Littell et al., 2017) which contains precomputed language distances based on the typological language information queried from the URIEL knowledge base  [^5] https://www.cs.cmu.edu/~dmortens/projects/7_project/. Following Ploeger et al. (2024), we calculate the distances between all pairs of languages in our dataset that have at least 5% coverage in the URIEL vectors. Figure 2 shows the distribution of pairwise geographic, syntactic, and genetic distances of all covered languages. We find that the languages in MultiQ cover a wide range of typologically similar and distant language pairs with an expected high skewness of genetic distance complementing previous research (Ploeger et al., 2024). We can therefore confidently speak of a high typological diversity of our dataset.
Figure 2: Distributions of the pairwise lang2vec distances for each language pair present in MultiQ.
Next, we calculate the typological feature coverage of the 137 languages in MultiQ, using language features provided in the Grambank database (Skirgård et al., 2023). For this purpose, we map the language IDs in MultiQ (obtained from the Google Translate API) to the Glottoids used by Grambank. [^6] For 14 languages in our dataset there is no matching entry in Grambank. To avoid incorrect mapping, we do not assign them manually and exclude these languages from the calculation Taken together, the languages in MultiQ cover at least 95.38% of the typological features recoded in Grambank. This underlines the typological diversity of our dataset.
Finally, we classify each language by its language family using the World Atlas of Language Structure (WALS).  [^7] https://wals.info/languoid In total, the 137 languages in MultiQ belong to 20 different language families. [^8] For languages that cannot be found in WALS (e.g.   Corsican), we manually look up the language family in Grambank. For additional details on the languages in our dataset, see Appendix A.
3 Experiments and Results
Using MultiQ, we can now answer our two main research questions regarding the multilingual language fidelity and multilingual QA accuracy of current chat-optimized open LLMs. We first separately assess fidelity and accuracy, and then evaluate their relationship. [^9] We publish all data and code at https://github.com/paul-rottger/multiq
3.1 Overall Experimental Setup
Models
We test six open-access LLMs that are both popular and competitive in performance with other state-of-the-art models as measured on standard (English-language) benchmarks such as the LMSys Leaderboard  [^10] https://huggingface.co/spaces/lmsys/chatbot-arena-leaderboard and AlpacaEval. [^11] https://tatsu-lab.github.io/alpaca_eval/ Specifically, the 7B, 13B and 70B versions of Llama2-Chat (Touvron et al., 2023b), the 7B Mistral-Instruct-v0.1 (Jiang et al., 2023), the 8x7B Mixtral-Instruct-v0.1 (Jiang et al., 2024) and the 7B Qwen1.5-Chat model (Bai et al., 2023). We test three sizes of Llama2 to evaluate scaling on MultiQ. Llama2 and Mistral are explicitly intended for English use only, whereas Mixtral and Qwen are explicitly multilingual: Mixtral “handles English, French, Italian, German and Spanish”, while Qwen offers unspecified “multilingual support”.
Inference
We run all models on two A100 GPUs using the simplegen Python library (Attanasio, 2023). We use default generation parameters from the transformers library, except for temperature, which we set to 0 to make completions deterministic. The maximum length of generations is 256 tokens. We do not use any system prompts. When prompting models with MultiQ questions, we do not provide any additional context or examples.
3.2 Language Fidelity
We gather responses from all six models described above on the 27,400 questions in MultiQ and then use GlotLID (Kargaran et al., 2023) to identify the response language. GlotLID is an open-source language identification model that supports more than 1,600 languages. GlotLID returns iso_636_9 language codes, which we manually map to the language codes in MultiQ. [^12] For 13 languages, several language codes in MultiQ map to just one iso_636_9 code. For details see Appendix B. Two languages in MultiQ, namely Meiteilon (Manipuri) and Dogri, are not supported by GlotLID, so we exclude them from our language fidelity analysis. Figure 3 shows high-level results on language fidelity, split by how often models responded in the language of the input prompt, in English, or another language.
We find that the Llama2 models show a very low language fidelity, responding predominantly in English, matching its intended use for English only. Fidelity increases with scale, but even Llama2 70b, which gives 21.4% answers in the prompt language, is much less faithful than the other models. Surprisingly, Mistral, also intended for English use only, shows the greatest language fidelity, giving 62.3% of answers in the prompt language. Mistral is closely followed by Mixtral (60.6%) and Qwen (50.4%), which are advertised as having multilingual capabilities. Interestingly, compared to Llama2, the other models more frequently opt neither for English nor the prompt language, but some other language in their response. This effect appears most evident for Qwen.
Figure 3: Overall language fidelity. Proportion of model responses (%) in the same language as the input prompt, in English, or in another language. We evaluate the responses of six models for 200 prompts in 135 languages (excl. Dogri & Meiteilon).
Next, we conduct a more fine-grained analysis of model answers in “another language”. For this, we focus on the three models responding in another language more than 10% of the time (Qwen, Mistral, and Mixtral). Of these, Mistral demonstrates the highest level of diversity among the languages it responds in. As opposed to Mixtral and Qwen, which draw from a pool of around 240 languages, GlotLID identifies more than 360 distinct languages in Mistral’s responses. We take a look into the correspondence between the language family of the input and output language. While Mixtral provides a response in the same language family at least 68.7% of the time, if not in an exact match or responding in English, Qwen demonstrates this behavior only 60.6% of the time, while Mistral does so for 54.6%. Note, however, that for 6% of the answers of Mistral and 7% of the answers of Mixtral, the language family of the response languages could not be determined and is thus classified as ’unknown’. Additionally, the distribution of language families in MultiQ is highly skewed, with 45% of the languages belonging to the ’Indo-European’ family, which spans a very broad range of languages from Irish to Turkish.
Figure 4: Granular language fidelity. Correlation matrices illustrating the relationship between input prompt and model response languages, shown as percentages. Axis ticks are selectively labeled for better visualization.
Finally, we analyze the relationship between the language of the prompt and the frequency of the respective response language of the models. Figure 4 shows correlation matrices for Mistral and Qwen. Corresponding matrices for the other models are shown in Appendix C. We find that Hindi is the most frequently selected language outside of the input language or English, closely followed by Indonesian. For Qwen, for example, these languages make up 20.4% of the “another language” category. This is visible by the thin yellow vertical lines leading to the axis ticks ind and hin in the Figure for both models respectively. It appears that the models lack support for numerous languages from India and Indonesia, thus treating Hindi and Indonesian as some kind of ’fallback’ languages for the wider language area. For example, the models often respond to languages such as Malay and Javanese in Indonesian, and languages such as Maithili, Konkani, and Bhojpuri in Hindi. However, this phenomenon can also be seen in smaller European languages. For example, the three models respond to questions in Croatian mostly in Italian, to those in Luxembourgish mostly in German, and those in Galician mostly in Portuguese. These observations underline the importance of improving multilingual models in the language coverage of low-resource languages.
Model Label P R F1
Mistral-7B Incorrect A. 1.00 1.00 1.00
Correct A. 1.00 1.00 1.00
Qwen1.5-7B Incorrect A. 0.89 0.99 0.94
Correct A. 0.95 0.67 0.78
Llama-2-7b Incorrect A. 0.90 1.00 0.95
Correct A. 1.00 0.70 0.83
Mixtral-8x7B Incorrect A. 0.85 0.98 0.91
Correct A. 0.97 0.76 0.85
Table 2: Evaluation of the GPT-4 classifier that we use to assess answer accuracy on MultiQ. Precision (P), Recall (R), and F1-Score of the GPT-4 judgments compared to human evaluation on answers from four different models to the same 100 random questions.
3.3 Question Answering Accuracy
Next, we assess how often models give correct answers, regardless of whether the language of this answer matches the input prompt or not.
Automated Evaluation
Since questions in MultiQ are open-ended and answers can come in many languages, we need a flexible method for evaluating accuracy. We find that a carefully crafted prompt to GPT-4, which checks a given model answer against the English version of the question from MultiQ, serves this purpose well. [^13] See Appendix B for the prompt template.
To evaluate the reliability of our automated evaluation method, we tasked two independent human annotators to label model responses from four of our models for the same 100 randomly selected prompts as correct or incorrect. [^14] We exclude the 13B and 70B versions of Llama2 from the QA accuracy analysis for reasons of clarity, and because they resemble the results of Llama2 7B. Disagreements between the annotators, which occurred for no more than four responses per model, were resolved in discussions with one of the authors. We find that the accuracy of our automated evaluation, as measured against the human labels, is very high across models (see Table 2). Accuracy is perfect on Mistral. For the other models, the automated evaluation tends to be very precise on correct answers, but less so on incorrect answers. This means that the automated evaluation will likely underestimate the proportion of correct model answers across languages. Overall, automated evaluation, like automated translation, introduces noise to our silver standard MultiQ benchmark, but we find the amount of noise to likely be small.
Results
Figure 5: Answer accuracy on MultiQ in proportion (%) of correctly answered questions per language. We compare four models of the same size across 137 languages and sort the results by median accuracy.
Model All en $\blacktriangle$10 $\blacktriangle$20 $\blacktriangle$50
Qwen (7B) 16.6 84.0 61.6 50.2 34.3
Mistral (7B) 15.4 84.5 64.6 56.6 37.6
Mixtral (8x7B) 33.8 90.5 87.7 84.8 67.4
Llama2 (7B) 19.2 81.5 58.4 53.9 41.4
Llama2 (13B) 23.4 82.0 66.4 62.7 49.6
Llama2 (70B) 29.1 90.5 80.7 76.5 61.2
Table 3: QA accuracy on MultiQ (%). We show accuracy overall, on English questions, and on the top($\blacktriangle$) 10, 20 and 50 best-performing languages for each model. Highest accuracy across models is bold.
Based on our automated evaluation, we calculate the proportion of correctly answered questions for each model across all 137 languages. Table 3 shows overall results, and Figure 5 shows a breakdown across all languages. We find that Mixtral is most accurate overall, but also across most individual languages covered in MultiQ. Mistral shows the lowest overall accuracy across all languages, with a long tail distribution starting to decrease after the best-performing 20 languages. Moreover, a direct comparison between the results of Mixtral and Qwen shows that they achieve very similar results on Chinese, although only Qwen was explicitly trained in Chinese (Bai et al., 2023).
Model Same English Other
Qwen (7B) 21.5 11.4 11.7
Mistral (7B) 17.1 14.6 11.8
Mixtral (8x7B) 35.0 37.0 26.7
Llama2 (7B) 44.7 14.9 46.5
Llama2 (13B) 51.6 15.1 49.6
Llama2 (70B) 61.4 17.3 48.2
Table 4: QA accuracy on MultiQ (%) split by language fidelity, i.e. which language models answered in (see Figure 3). Highest accuracy per model is bold.
3.4 Language Fidelity vs Answer Accuracy
Finally, we combine the results of §3.2 and §3.3 to assess the relationship between language fidelity and QA accuracy. Table 4 shows the mean accuracy grouped by response language category. We find that Llama2, despite its overall low language fidelity, shows strong QA accuracy. Especially when answering in the same language as prompted, it gives a correct answer in almost half of the cases (44.7%). Scaling the model further improves accuracy, with the 13B and 70B variants achieving higher correct answer rates of 51.6% and 61.4%, respectively, when answering in the same language as prompted. Qwen and Mistral show a similar pattern, i.e.  higher answering accuracy when answering in the prompt language with 21.5% and 17.1%. However, their overall accuracy is quite low, especially when compared to their initial high language fidelity. Strikingly, we find significant differences across models that ought to be similar due to their intended use for English, i.e.  Llama2 demonstrating low fidelity yet high accuracy, in contrast to Mistral, which exhibits the reverse pattern. A qualitative analysis of Mistral’s answers reveals that the model often merely repeats the questions it was asked, which is technically faithful but never an accurate answer. Only for Mixtral is this pattern not apparent, with the highest accuracy of 37% being achieved for responses in English. Overall, models answering in the prompted language tend to be more accurate than those responding in English, Mixtral being the exception. Therefore, our results suggest that increased language fidelity may positively impact QA accuracy.
4 Drivers of Multilinguality
Our results show significant variations in terms of language fidelity, QA accuracy, and the relationship between them, across the models we test with MultiQ. This prompts us to investigate what factors may explain these differences. We focus on the differences between Mistral and Llama2 7B, which are particularly surprising given that both models are intended for use in English only. Prior research highlights the significant roles of tokenization and training data in multilingual capability (Dufter and Schütze, 2020; Clark et al., 2022; Petrov et al., 2023). Since there is little to no public information on the training data of the models that we test, we focus on their tokenizers.
Background: Byte Pair Encoding
Both Llama2 and Mistral employ a tokenizer that uses byte pair encoding (BPE) (Sennrich et al., 2016), streamlining sequence encoding by minimizing token count through identifying common subwords. This allows for frequent terms to be represented through subwords, whereas rare or unseen terms prompt the model to default to individual characters or ultimately ASCII codes for representation.
Unique Tokens in MultiQ
First, we evaluate how different tokenization strategies impact prompt representation in MultiQ. We observe a significant deviation in the number of unique tokens used to represent all prompts in MultiQ: Mistral uses 9,933 unique tokens, contrasting with Llama2’s 10,676. This suggests that Llama2 may be less efficient at segmenting the typologically diverse MultiQ dataset into a smaller number of subwords.
Tokenization Strategies
For each model, we group languages into three categories depending on the model’s tokenization strategy fo the respective language. We develop a heuristic that allows us to classify languages into three tokenization categories: “ASCII”, “character” or “subword”. We do so by compiling the model’s 20 most commonly used tokens for representing each language in MultiQ, removing noise tokens (e.g., sentence start) from the list, and then quantifying the prevalence of ASCII tokens and characters based on the token-id ranges they usually occupy. Languages with over 70% of tokens in the ASCII or character categories were classified accordingly. For Llama2, out of 137 languages, we find 89 subword, 36 character and 12 ASCII languages. Mistral has 37 in character and 9 in ASCII. The models differ in the tokenization of some symbolic languages such as Chinese (zh) (see Figure 6), compared to Mistral mostly using individual characters, Llama2 more frequently has to use ASCII tokens due to its limited recognition of Chinese tokens, hindering its ability to effectively map Chinese vocabulary.
Tokenization vs. QA Accuracy
We evaluate the models’ average QA accuracy across tokenization categories, finding a clear hierarchy (see Table 5): Subword encoding outperforms character and ASCII encodings, with Mistral achieving 20.4% accuracy on subword encoded languages and Llama2 24%. Acknowledging the limitations of our heuristic in precisely classifying tokenization strategies, we believe that exploring tokenization optimization is a promising direction for multilingual research, which also aligns with the broader discourse on the impact of tokenization on model capabilities. We hope these insights highlight the importance of alternative language representation strategies (e.g., pixel-based models (Salesky et al., 2023)).
Figure 6: Tokenization analysis. Distribution of the most used unique tokens to represent all prompts in Chinese (zh), given the respective tokenizers of the Llama2 (7B) and Mistral (7B).
Model Subword Character ASCII
Llama (7B) 24.0 11.15 8.0
Mistral (7B) 20.4 6.4 2.1
Table 5: QA accuracy on MultiQ split by tokenization strategy. Highest accuracy per model is bold.
5 Related Work
We discuss the related literature with respect to (i) large multilingual benchmarks, (ii) explicitly multilingual models, and (iii) multilingual studies of monolingual models.
Multilingual Benchmarks.
Existing multilingual benchmarks primarily target the performance of fully or partially supervised models on collections of standard NLP tasks, like XNLI (Conneau et al., 2018). Popular examples are the XTREME benchmarks (e.g., Ruder et al., 2021, 2023), and XGLUE (Liang et al., 2020). Similarly, researchers also presented benchmarks designed for languages spoken in particular regions, e.g., IndicXTREME (Doddapaneni et al., 2023) for Indic languages, Masakhan-NER (Adelani et al., 2022), and TaTA (Gehrmann et al., 2023) covering African languages, as well as NusaX (Winata et al., 2023) for Indonesian languages. Recently, Ahuja et al. (2023) proposed MEGA for evaluating multilingual generative models, which they use to evaluate LLMs like GPT-4 on a set of standard tasks. In a similar vein, Asai et al. (2023) presented BUFFET and benchmark LLMs for few-shot transfer. Artetxe et al. (2020) present XQuaD, a multilingual QA dataset focusing on reading comprehension. All these benchmarks typically cover roughly 20 to 90 languages. In contrast, MultiQ contains parallel questions in 137 languages covering 95.38% of Grambank features, which allows us to study the multilingual behavior of LLMs at the margins.
Multilingual LLMs.
In the past, researchers have presented a range of multilingual LMs like mBERT (Devlin et al., 2019), XLM-R (Conneau et al., 2020), and Glot-500 ImaniGooghari et al. (2023). These models are most effective in a supervised fine-tuning setting. Different from that, models like mT5 (Xue et al., 2021), BLOOMZ (Workshop et al., 2022), Okapi (Lai et al., 2023b) and the very recently released Aya (Üstün et al., 2024) better allow for prompting. For all those models, many studies also explored how to further adapt them to more languages (e.g., Yong et al., 2023a), assessed their detailed capabilities (e.g., Zhang et al., 2023a; Yong et al., 2023b), and aspired to understand the cross-lingual mechanisms underpinning their capabilities (e.g., Lauscher et al., 2020; Choenni et al., 2023, inter alia). Given the scale of large chat models, the landscape of open models that cover multiple languages is, however, still scarce.
Multilinguality in Monolingual LLMs.
Given that multilingual chat models are barely openly available, we are especially interested in assessing the multilinguality of supposedly English-only models. Blevins and Zettlemoyer (2022) explained this behavior through data contamination: while the vast majority of the pre-training data of those models is English (e.g., ~93% for GPT-3 (Brown et al., 2020), and ~90% for Llama-2 (Touvron et al., 2023a)) there are also small portions of non-English training data present. In such cases, the dominant language can help unlock the models’ capabilities for the underrepresented languages (Gogoulou et al., 2022). As a result, recent research assesses the multilinguality of several English-centric models, like GPT-3 and ChatGPT (e.g., Zhang et al., 2023b; Lai et al., 2023a; Armengol-Estapé et al., 2022; Winata et al., 2021). However, all these works focus on standard NLP tasks, one or few models, and on a few major languages only, while we analyze the multilingual behavior of six LLMs at scale.
6 Conclusion
We introduced MultiQ, a new silver standard benchmark for open-ended question answering that covers 137 typologically diverse languages. With MultiQ, we evaluated the basic multilingual capabilities of six current, chat-optimized open LLMs, which are not all intended for multilingual use. Our analysis focused on two key dimensions of multilingual capability – language fidelity and QA accuracy – and how they relate to each other. We found that some models designed primarily for English reveal significant capabilities in one dimension or the other. We noticed that although language fidelity is not decisive, it positively impacts QA accuracy, prompting us to explore other underlying factors. We hope that our work will spark further research in the direction of the multilingual capability of open chat-optimized models, with a particular emphasis on diverse multilingual dimensions.
Limitations
Our work comes with several limitations that have already been partially discussed throughout the paper. First, we automatically translate our dataset, which introduces noise in the dataset, however, given our detailed validation, we strongly believe that the implications of our findings still hold. Second, the automated evaluation of the models’ answering accuracy using GPT-4 and their response language using GlotLID are additional factors that might bring noise into the results. However, further validation assures us that these factors minimally impact the results, potentially leading to an underestimation of answer accuracy. Third, mapping Google Translate IDs to the ISO language codes of GlotLID wasn’t always directly possible, but we excluded possible inaccuracies from our calculations. Lastly, our analysis intentionally concentrated on basic multilingual capabilities, excluding the assessment of advanced reasoning or formulation skills.
Ethical Considerations
Intended Use
As we emphasized throughout our paper, MultiQ is intended to test basic multilingual capabilities. Therefore, good performance on MultiQ alone should not be used as evidence for an LLM being suitable for specific languages.
Acknowledgements
The work of Carolin Holtermann and Anne Lauscher is funded under the Excellence Strategy of the German Federal Government and the States. Paul Röttger is a member of the Data and Marketing Insights research unit of the Bocconi Institute for Data Science and Analysis, and is supported by a MUR FARE 2020 initiative under grant agreement Prot. R20YSMBZ8S (INDOMITA). We thank our native-language annotators for their valuable work.
Appendix
Appendix A MultiQ
In total MultiQ covers 10 different question domains and 20 distinct language families.
Domains
  • chemistry
  • physics
  • astronomy
  • history
  • maths
  • geography
  • art
  • sports
  • music
  • animals
Language Families
  • Afro-Asiatic (AA)
  • Altaic (Al)
  • Austro-Asiatic (AuA)
  • Austronesian (Au)
  • Aymaran (Ay)
  • Basque (B)
  • Dravidian (D)
  • Hmong-Mien (HM)
  • Indo-European (IE)
  • Japanese (J)
  • Kartvelian (K)
  • Korean (Ko)
  • Mande (M)
  • Niger-Congo (NC)
  • Other (O)
  • Quechuan (Qu)
  • Sino-Tibetan (ST)
  • Tai-Kadai (TK)
  • Tupian (T)
  • Uralic (U)
Language Prop. of correct questions
Chinese 0.955
Farsi 0.890
French 0.915
German 0.995
Indonesian 0.889
Italian 0.990
Japanese 0.955
Spanish 0.950
Urdu 0.935
Table 6: Validation Results on MultiQ. We present the proportion of correctly translated prompts for each language assessed by native speakers in the respective language.
Code
Language
Family Code
Language
Family Code
Language
Family
af
Afrikaans
IE hmn
Hmong
HM ny
Nyanja (Chichewa)
NC
ak
Twi (Akan)
NC hr
Croatian
IE om
Oromo
AA
am
Amharic
AA ht
Haitian Creole
or
Odia (Oriya)
IE
ar
Arabic
AA hu
Hungarian
U pa
Panjabi
IE
as
Assamese
IE hy
Armenian
IE pl
Polish
IE
ay
Aymara
Ay id
Indonesian
Au ps
Pashto
IE
az
Azerbaijani
Al ig
Igbo
NC pt
Portuguese
IE
be
Belarusian
IE ilo
Ilocano
Au qu
Quechua
Qu
bg
Bulgarian
IE is
Icelandic
IE ro
Romanian
IE
bho
Bhojpuri
IE it
Italian
IE ru
Russian
IE
bm
Bambara
M iw
Hebrew alternativ
AA rw
Kinyarwanda
NC
bn
Bengali
IE ja
Japanese
J sa
Sanskrit
bs
Bosnian
IE jv
Javanese alternativ
Au sd
Sindhi
IE
ca
Catalan
IE jw
Javanese
Au si
Sinhala (Sinhalese)
IE
ceb
Cebuano
Au ka
Georgian
K sk
Slovak
IE
ckb
Kurdish (Sorani)
IE $k$
Kazakh
Al sl
Slovenian
IE
co
Corsican
IE km
Khmer
AuA sm
Samoan
Au
cs
Czech
IE kn
Kannada
D sn
Shona
NC
cy
Welsh
IE ko
Korean
K so
Somali
AA
da
Danish
IE kri
Krio
sq
Albanian
IE
de
German
IE ku
Kurdish
IE sr
Serbian
IE
doi
Dogri
IE ky
Kyrgyz
Al st
Sesotho
NC
dv
Dhivehi
IE la
Latin
su
Sundanese
Au
$e$
Ewe
NC lb
Luxembourgish
IE sv
Swedish
IE
el
Greek
IE lg
Luganda
NC sw
Swahili
NC
en
English
IE ln
Lingala
NC ta
Tamil
D
eo
Esperanto
IE lo
Lao
TK te
Telugu
D
es
Spanish
IE lt
Lithuanian
IE tg
Tajik
IE
et
Estonian
U lus
Mizo
ST th
Thai
TK
eu
Basque
B lv
Latvian
IE ti
Tigrinya
AA
fa
Persian
IE mai
Maithili
IE tk
Turkmen
Al
fi
Finnish
U mg
Malagasy
Au tl
Tagalog (Filipino)
Au
fil
Filipino (Tagalog)
Au mi
Maori
Au tr
Turkish
Al
fr
French
IE mk
Macedonian
IE ts
Tsonga
NC
fy
Frisian
IE ml
Malayalam
D $t$
Tatar
Al
ga
Irish
IE mn
Mongolian
ug
Uyghur
Al
gd
Scots Gaelic
IE mni-Mtei
Meiteilon (Manipuri)
ST uk
Ukrainian
IE
gl
Galician
IE mr
Marathi
IE ur
Urdu
IE
gn
Guarani
T ms
Malay
Au uz
Uzbek
Al
gom
Konkani
IE mt
Maltese
AA vi
Vietnamese
AuA
gu
Gujarati
IE my
Myanmar (Burmese)
ST xh
Xhosa
NC
ha
Hausa
AA ne
Nepali
IE yi
Yiddish
IE
haw
Hawaiian
Au nl
Dutch
IE yo
Yoruba
NC
he
Hebrew
AA no
Norwegian
IE zh
Chinese (Trad.)
hi
Hindi
IE nso
Sepedi
NC zh-CN
Chinese (Simpl.)
zh-TW
Chinese (Simpl.)
zu
Zulu
NC
Table 7: 137 Languages covered by MultiQ, we present their Google Translate Code and the acronym of their language families.
Appendix B Experimental Setup
In Table 8, we provide all information and links to the data, models, and frameworks, we use in our work. All artifacts were used according to their intended use, as described in their licenses. As described in the main body of this manuscript, we are also releasing our code publicly and will do so under the CC-BY-4.0 license.
Purpose Name Vocabsize Intended Languages
Models meta-llama/Llama-2-7b-chat-hf 32000 en
meta-llama/Llama-2-13b-chat-hf 32000 en
mistralai/Mistral-7B-Instruct-v0.1 32000 en
Qwen/Qwen1.5-7B-Chat 151936 undefined
mistralai/Mixtral-8x7B-Instruct-v0.1 32000 fr, en, es, it, de
cis-lmu/glotlid 1600 langs
gpt-4-0125-preview
Framworks transformers==4.34.0
torch==2.1.0
huggingface-hub==0.17.3
tokenizers==0.14.1
Datasets WALS           Dataset License: CC-BY-4.0
Grambank           License CC-BY-4.0
Table 8: Overview of models, datasets and frameworks used in our work.
Task
Prompt
Question Creation for MultiQ
Generate a list of 10 simple and clear {domain} questions with simple and clear answers.
Validate Answer Correctness
For the following question, is this a correct answer?
Question: {MultiQ Question in English}
Answer: {Model Answer}
The question is in English. The answer is either in English or in another language. If the answer is correct, type ’yes’. If the answer is incorrect, type ’no’."
Table 9: Prompt templates used to prompt GPT-4 for question creation and QA accuracy validation respectively.
Google Translate ID ISO_639_3
zh-CN zho
zh
zh-TW
ku ckb
ckb
he heb
iw
sr hbs
hr
jv jav
jw
tl tgl
fil
Table 10: List of Languages in MultiQ that map to only one iso code in the GlotLID model.
Appendix C Granular Language Fidelity Analysis
Figure 7: Correlation matrices illustrating the relationship between input question languages and model response languages for the Llama2 models, shown as percentages. Axis ticks are selectively labeled for clearer visualization.
Figure 8: Correlation matrix illustrating the relationship between input question languages and model response languages for Mistral, shown as percentages. Axis ticks are selectively labeled for clearer visualization.
Appendix D Answer Accuracy
Figure 9: Answer Accuracy on MultiQ in proportion (%) of correctly answered questions per language. We compare the Llama2 models in different model sizes across all 137 languages and sort the results by median accuracy.